OpenStudy (anonymous):

Simplify: (5x2 + 3x + 4) − (5x2 + 5x − 1) 8x + 5 −2x + 5 10x2 + 8x + 3 10x2 − 2x + 3

OpenStudy (anonymous):

What do you think? Have you tried??

OpenStudy (anonymous):

\[(5x^2 + 3x + 4) − (5x^2 + 5x − 1)\]

OpenStudy (anonymous):

I got -2x+5

OpenStudy (anonymous):

Firstly you should know : -(+) = - -(-) = +

OpenStudy (skullpatrol):

Subtracting polynomials is like subtracting real numbers. To subtract a real number you add the opposite of that number. To subtract a polynomial you add the opposite of each term of that polynomial, and then simplify.

OpenStudy (anonymous):

Ok

OpenStudy (skullpatrol):

your answer is right @kaylamarie081 :)

OpenStudy (anonymous):

We will open second brackets by multiplying -ve sign with it.. So: \(-(5x^2 + 5x - 1) = -5x^2 - 5x + 1\)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You should note how +5x changes to -5x and how -1 changes to +1..

OpenStudy (skullpatrol):

Thanks for asking :)

OpenStudy (skullpatrol):

$$\Huge -(a+b)=(-a)+(-b)$$

OpenStudy (skullpatrol):

To find the opposite of a sum or difference you change the sign of each term of the sum or difference.

OpenStudy (skullpatrol):

This is just a special case of the distributive property with a = -1. $$-1(a+b)=-1a+(-1)b$$

OpenStudy (skullpatrol):

where we know -1a = -a